### v1

@@ 5 Related Work
Epistemic Uncertainty Collapse. Recent work by Fellaji & Pennerath (2024) observed decreasing epistemic uncertainty as model size and dataset size vary, even in simple MLPs. They termed the decrease in epistemic uncertainty the “epistemic uncertainty hole” but left its explanation to future work. Our study provides an explanation through both a theoretical framework and additional empirical evidence via ensembles of ensembles, implicit ensembling, and ensemble extraction. A more detailed comparison can be found in Appendix C.
Extracting Sub-Models from a Larger Network. The concept of extracting sub-models from a larger network shares similarities with several existing approaches in the literature. The “lottery ticket hypothesis” (Frankle & Carbin, 2018) proposes that dense, randomly-initialized networks contain sparse subnetworks capable of training to similar accuracy. However, our approach differs in that we do not retrain the subnetworks, but rather identify diverse substructures within the pre-trained model. Our method is more closely related to “sub-network ensembles” (Durasov et al., 2021), where multiple subnetworks are extracted from a single trained network to form an ensemble. Unlike previous work that primarily focused on pruning for efficiency, our approach aims to recover epistemic uncertainty. We introduce a novel mutual information-based objective to obtain diverse masks, emphasizing diversity rather than pruning. This allows us to extract an ensemble that better captures the model’s internal epistemic uncertainty.
@@
In contrast to these works, this work addresses fundamentally different research questions. We provide a novel, unified explanation for epistemic uncertainty collapse in large models and hierarchical ensembles, a phenomenon not previously explored in depth. Furthermore, we introduce and examine implicit ensemble extraction to mitigate epistemic uncertainty collapse.
@@ 6 Conclusion
This work has uncovered and analyzed the collapse of epistemic uncertainty in large neural networks and hierarchical ensembles. Our theoretical framework and empirical results demonstrate this phenomenon across various architectures and datasets, from simple MLPs to state-of-the-art vision models.
To address this issue, we examined implicit ensemble extraction, a method to decompose larger models into sub-models. This approach effectively recovers hidden ensemble structures and improves epistemic uncertainty estimates, supporting our hypothesis that implicit ensembling contributes to the observed uncertainty collapse.
Our findings challenge the assumption that more complex models invariably offer better uncertainty quantification out of the box. They open new avenues for research into uncertainty estimation in large-scale machine learning models, particularly for safety-critical applications and out-of-distribution detection tasks. Future work should focus on developing robust methods to maintain reliable uncertainty estimates as model complexity increases, ensuring the continued advancement of trustworthy AI systems.
@@ Acknowledgments
We thank Jannik Kossen and Tom Rainforth for early discussions after an initial presentation at an OATML research talks. The model checkpoints used for the initial explicit ensemble of ensemble experiments on CIFAR-10 were trained by Jishnu Mukhoti as part of Mukhoti et al. (2023) and re-used here. We also thank Alex Evans for feedback on an earlier draft, and Midjourney for generously supporting this personal research.
Appendix A Model and Dataset Details
A.1 MNIST Experiments
@@ A.3 ImageNet Experiments
A.4 Implicit Ensemble Extraction
@@
Appendix B Evaluation Details
@@
B.1 Weighted Mutual Information
@@ B.2 Calibration Error
B.3 Metric Computation by Uncertainty Score
@@
B.4 Other Evaluation Metrics
@@
Appendix C Comparison with Fellaji & Pennerath (2024)
Our work explains the main findings of Fellaji & Pennerath (2024), who observe what they term the “epistemic uncertainty hole” in Bayesian neural networks. While their work also provides empirical evidence for this phenomenon, our study offers several key extensions and insights:
Theoretical Framework: We provide a theoretical explanation for the epistemic uncertainty collapse through the lens of ensembles of ensembles and implicit ensembling, offering a mechanistic understanding of why this phenomenon occurs:
Ensembles of Ensembles: Our work introduces the concept of ensembles of ensembles, showing how the epistemic uncertainty collapse manifests in hierarchical ensemble structures. This provides a novel perspective on the phenomenon not explored in the original work.
@@ Appendix C Comparison with Fellaji & Pennerath (2024)
Broader Model Architectures: While Fellaji & Pennerath (2024) primarily focused on MLPs, we demonstrate that this phenomenon extends to more complex architectures, including state-of-the-art vision models based on ResNets and Vision Transformers.
Thus, while Fellaji & Pennerath (2024) observe an epistemic uncertainty collapse, our work provides a more comprehensive theoretical and empirical investigation of this phenomenon. We not only confirm their findings across a broader range of models and datasets but also offer new insights into the mechanisms behind this effect and potential strategies for mitigation.
Furthermore, our analysis of the implications for out-of-distribution detection and the proposed implicit ensemble extraction technique represent initial steps towards addressing the practical challenges posed by the epistemic uncertainty collapse in real-world applications.
Appendix D Additional Results

### v2

@@ 6 Related Work
Epistemic Uncertainty Collapse. A recent preprint by Fellaji & Pennerath (2024) observed decreasing epistemic uncertainty as model size and dataset size vary, even in simple MLPs. They termed the decrease in epistemic uncertainty the “epistemic uncertainty hole” but left its explanation to future work. Our study proposes an explanation through a theoretical framework for ensembles of ensembles, the hypothesis of implicit ensembling for wider models, and additional empirical evidence from experiments on ensemble extraction. A more detailed comparison can be found in Appendix D.
Extracting Sub-Models from a Larger Network. The concept of extracting sub-models from a larger network shares similarities with several existing approaches in the literature. The “lottery ticket hypothesis” (Frankle & Carbin, 2018) proposes that dense, randomly-initialized networks contain sparse subnetworks capable of training to similar accuracy. However, our approach differs in that we do not retrain the subnetworks, but rather identify diverse substructures within the pre-trained model. Our method is more closely related to “sub-network ensembles” (Durasov et al., 2021), where multiple subnetworks are extracted from a single trained network to form an ensemble. Unlike previous work that primarily focused on pruning for efficiency, our approach aims to recover epistemic uncertainty. We introduce a novel mutual information-based objective to obtain diverse masks, emphasizing diversity rather than pruning. This allows us to extract an ensemble that better captures the model’s internal epistemic uncertainty. This goal is more similar to Wortsman et al. (2020) which is concerned with discovering sub-networks during training for continual learning of diverse tasks without catastrophic forgetting. Our approach is only applied post-training, however.
@@
In contrast to these works, this work addresses fundamentally different research questions. We propose a novel, unified explanation for epistemic uncertainty collapse in large models and hierarchical ensembles, a phenomenon not previously explored in depth. Furthermore, we introduce and examine implicit ensemble extraction to mitigate epistemic uncertainty collapse.
Relationship to Feature and Neural Collapse. Our findings on epistemic uncertainty collapse share interesting connections with, but remain distinct from, two related phenomena in deep learning. Feature collapse, introduced by Van Amersfoort et al. (2020), describes how intermediate feature representations become increasingly similar during training, leading to overconfident predictions. Neural collapse (Papyan et al., 2020) characterizes how features and classifiers converge to a specific geometric structure in the terminal phase of training.
While these phenomena involve forms of representation collapse, our observed epistemic uncertainty collapse operates at a different level. Rather than focusing on feature geometry or representation similarity, we study how predictive uncertainty estimates degrade in two key settings: (1) ensembles of ensembles, where individual ensembles converge to similar predictive distributions despite diverse training, and (2) large models through the proposed implicit ensembling, where we can extract diverse sub-models that reveal hidden uncertainty structure. The distinction becomes particularly clear in regression tasks (Figure 1), where neural collapse theory does not apply due to its reliance on classification-specific geometric properties.
@@ 7 Conclusion
This work has analyzed the collapse of epistemic uncertainty in large neural networks and hierarchical ensembles. Our theoretical framework and empirical results demonstrate this phenomenon across various architectures and datasets, from simple MLPs to state-of-the-art vision models.
To further explore this phenomenon, we examined implicit ensemble extraction, a method to decompose larger models into sub-models. This approach can recover diverse predictive distributions from a single model and improve epistemic uncertainty estimates. The effectiveness of this extraction method is consistent with our implicit ensembling hypothesis, though it does not definitively establish it as the underlying mechanism of epistemic uncertainty collapse.
Our findings challenge the assumption that more complex models invariably offer better uncertainty quantification out of the box. They open new avenues for research into uncertainty estimation in large-scale machine learning models, particularly for safety-critical applications and out-of-distribution detection tasks. Future work should investigate the exact mechanisms behind epistemic uncertainty collapse in large models and develop robust methods to maintain reliable uncertainty estimates as model complexity increases, ensuring the continued advancement of trustworthy AI systems.
@@ Acknowledgments
We thank the anonymous TMLR reviewers and Freddie Bickford Smith for their helpful feedback. Alex Evans provided valuable comments on an earlier draft, and Midjourney generously supported this personal research last summer. Jishnu Mukhoti trained the model checkpoints used in the initial explicit ensemble of ensemble experiments on CIFAR-10 as part of Mukhoti et al. (2023). Discussions with Lisa Schut helped shape the intuitions about NTK and implicit ensembling. We also appreciate Jannik Kossen and Tom Rainforth’s insights following an early presentation at OATML.
Appendix A Theoretical Framework
A.1 Variance-Based Epistemic Uncertainty
A convenient way to obtain a tractable upper bound for the mutual information uses the law of total variance on the random variable . Recall the law of total variance:
It is sometimes used as an uncertainty decomposition in its own right (Kendall & Gal, 2017), where the first term on the right is the average conditional variance across models, considered the aleatoric part, while the second term quantifies how the conditional mean changes with , considered the epistemic part.
To connect this decomposition to , assume each conditional distribution is (approximately) Gaussian with constant covariance, and approximate using a Gaussian as maximum entropy distribution. The differential entropy of a Gaussian random variable with variance is . This allows us to approximate by comparing the entropy of these Gaussians:
The numerator reflects epistemic variability (how much the mean shifts between models), and the denominator captures the typical conditional variance. This ratio then appears within a logarithm because differential entropy grows with the log of variance in the Gaussian setting.
From this, we can see that variance-based epistemic uncertainty can be loosely matched to mutual information when each conditional distribution is close to Gaussian with roughly constant variance, making total variance reflect the same factors that drive the mutual information. Outside this regime—particularly with multimodal, heavy-tailed, or strongly heteroscedastic predictive distributions—the variance-based approach masks richer dependencies in , and the two metrics diverge as mutual information accounts for more complex, higher-order effects222This section is based on an informal note from 2022: https://blackhc.notion.site/Law-of-Total-Variance-vs-Information-Theory-1608b02974c54839a7c4a8e181f75e7f..
A.2 Deep Ensembles as Approximate Posterior Sampling
The theoretical justification for using deep ensembles to estimate epistemic uncertainty stems from their connection to Bayesian inference (Wilson & Izmailov, 2020). When training neural networks with different random initializations and stochastic optimization, the resulting ensemble can be viewed as drawing approximate samples from the posterior distribution over model parameters (Stephan et al., 2017; Shen et al., 2024b).
This interpretation arises because the training process combines:
Random initialization, which provides diverse starting points in parameter space (Garipov et al., 2018)
Stochastic gradient descent, which introduces noise in optimization trajectories (Stephan et al., 2017)
Weight decay regularization, which acts as an implicit prior over parameters (Loshchilov & Hutter, 2019)
While deep ensembles do not perform exact Bayesian inference, they capture important aspects of posterior uncertainty through the diversity of solutions found (Fort et al., 2019). The independent training runs explore different modes of the loss landscape (Garipov et al., 2018), approximating the kind of parameter diversity we would expect from true posterior samples. This provides a practical foundation for using ensembles to estimate epistemic uncertainty, even without explicit variational approximations or MCMC sampling (Wilson & Izmailov, 2020).
The quality of this approximation improves with ensemble size, as more independent training runs better characterize the range of plausible models given the data (Lakshminarayanan et al., 2017). This theoretical connection motivates why deep ensembles often provide better uncertainty estimates than methods that make stronger assumptions about the form of the approximate posterior (Ovadia et al., 2019).
A.3 Epistemic Uncertainty Collapse in Ensembles of Ensembles
The rather informal exposition in Section 3.2 does not differentiate between empirical mutual information which measures the disagreement between the predictions of the ensemble members and the mutual information between the predictions and the parameter distribution where the ensemble members are seen as samples from . It is not necessary to understand the effect and the empirical results. Let us make this distinction clearer here for completeness’ sake.
The empirical epistemic uncertainty of a specific sub-ensemble is . The epistemic uncertainty of the paramter distribution itself is . We have:
where follows a uniform distribution. We can write this more explicitly as:
The entropy is a continuous function, so we can take the limit as the the number of models in the sub-ensemble goes to infinity and swap the order of limit and entropy, where we use that following (9) and similarly . That is, for , we have:
Now, crucially, we can make the same argument for the epistemic uncertainty of the ensemble comprising all models in all the sub-ensembles , and we obtain the same result:
With the chain rule of mutual information(11) we obtain:
and in the limit of infinitely-sized sub-ensembles, the epistemic uncertainty across the sub-ensembles is thus sandwiched to 0:
A.4 Epistemic Uncertainty Collapse in Random Forests
To further validate the epistemic uncertainty collapse phenomenon, we examine its manifestation in random forests—another ensemble learning method with well-understood theoretical properties. An ensemble of random forests (Breiman, 2001) naturally forms an ensemble of ensembles, where each random forest is an ensemble of decision trees.
We replicate the toy regression problem from earlier, training random forest ensembles with varying numbers of trees (1, 2, 4, 8, 16, 32, and 64) and creating ensembles of ensembles of 10 random forests for each configuration. Each random forest uses decision trees with a maximum depth of 3 to prevent overfitting while maintaining sufficient model capacity.
The results in Figure 6 mirror our findings with neural networks. As the number of trees in each random forest increases:
the ensemble predictions converge more tightly to the true function, with narrowing uncertainty bands;
the epistemic uncertainty decreases systematically across the input space, particularly in regions between training points; and
the mean epistemic uncertainty shows a clear inverse relationship with forest size
This parallel between random forests and neural networks underscores that epistemic uncertainty collapse is a general phenomenon in ensemble methods, rather than being specific to neural networks. The observation that this effect appears in random forests—which have stronger theoretical guarantees and more interpretable behavior than neural networks—provides additional evidence for the fundamental nature of this phenomenon.
Appendix B Model and Dataset Details
B.1 MNIST Experiments
@@ B.3 ImageNet Experiments
B.4 Implicit Ensemble Extraction
@@
Appendix C Evaluation Details
@@
C.1 Weighted Mutual Information
@@ C.2 Calibration Error
C.3 Metric Computation by Uncertainty Score
@@
C.4 Other Evaluation Metrics
@@
Appendix D Comparison with Fellaji & Pennerath (2024)
While we initially observed and documented the epistemic uncertainty collapse phenomenon in 2021, Fellaji & Pennerath (2024) independently discovered similar effects in Bayesian neural networks, terming it the “epistemic uncertainty hole”. Our study offers several key extensions and insights:
Theoretical Framework: We provide a theoretical explanation for the epistemic uncertainty collapse through the lens of ensembles of ensembles and propose the hypothesis of implicit ensembling for wider models, offering a mechanistic understanding of why this phenomenon occurs:
Ensembles of Ensembles: Our work introduces the concept of ensembles of ensembles, showing how the epistemic uncertainty collapse manifests in hierarchical ensemble structures. This provides a novel toy setting for the phenomenon not explored in the other work.
@@ Appendix D Comparison with Fellaji & Pennerath (2024)
Broader Model Architectures: While Fellaji & Pennerath (2024) primarily focus on MLPs, we demonstrate that this phenomenon extends to more complex architectures, including state-of-the-art vision models based on ResNets and Vision Transformers.
Thus, while Fellaji & Pennerath (2024) observe an epistemic uncertainty collapse, our work provides a more comprehensive theoretical and empirical investigation of this phenomenon. We not only confirm their findings across a broader range of models and datasets but also offer new insights into the mechanisms behind this effect and potential strategies for mitigation. The proposed implicit ensemble extraction represents an initial step towards addressing the practical challenges posed by the epistemic uncertainty collapse in real-world applications.
Appendix E Additional Results